# Babel-ImageNet: Massively Multilingual Evaluation of Vision-and-Language Representations

### v1

@@ 7 Conclusion
We introduced Babel-ImageNet, the first massively multilingual translation of the ImageNet-1k classes to 92 languages. We leverage the WordNet synsets as the link between ImageNet and BabelNet to obtain high-quality translations without relying on MT or human annotators. Using Babel-ImageNet, we carried out the most comprehensively multilingual comparative evaluation of 8 publicly available CLIP models on zero-shot image classification, demonstrating that all models fail for low(er)-resource languages. Crucially, we validate our benchmark by showing that models’ text-to-image retrieval performance (on three datasets) strongly correlates with their ZS-IC performance on Babel-ImageNet for the corresponding languages. Finally, we proposed a parameter-efficient fine-tuning procedure that drastically improves the performance of multilingual CLIP models for low-resource languages.
@@
Acknowledgments and Disclosure of Funding
This work was partly supported by the Alexander von Humboldt Foundation.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
  2. 2.
    Did you describe the limitations of your work? [Yes]
We analyze the limitations of our partial label translation in §5 and discuss our training limitations from §6.
Did you discuss any potential negative societal impacts of your work? [Yes] We discuss some potential issues in §B. Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] If you are including theoretical results…
  1. 1.
    Did you state the full set of assumptions of all theoretical results? [N/A]
  2. 2.
    Did you include complete proofs of all theoretical results? [N/A]
If you ran experiments…
  1. 1.
    Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
See §A Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See §C.3 Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] Yes where applicable but most results are single-seed (because zero-shot). Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Yes, but we report training time and GPUs only for §6; other sections are evaluation-only. If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
  1. 1.
    If your work uses existing assets, did you cite the creators? [Yes]
Did you mention the license of the assets? [No] See §A for licensing of our assets. Other licenses are not explicitly stated. Did you include any new assets either in the supplemental material or as a URL? [Yes] See §A Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] If you used crowdsourcing or conducted research with human subjects…
  1. 1.
    Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
  2. 2.
    Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
  3. 3.
    Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A License and Hosting
Babel-ImageNet is a processed version of BabelNet v5.2 downloaded from https://babelnet.org, made available with the BabelNet Non-Commercial License (see https://babelnet.org/full-license). All data, code to generate the data (both Babel-ImageNet and MT prompts), and evaluation code can be found at https://github.com/gregor-ge/Babel-ImageNet. We bear all responsibility in case of violation of rights.
Appendix B Limitations and Societal Impact
@@
While Babel-ImageNet greatly improves language coverage for the evaluation of multilingual vision-language models, there are some limitations of our work: For one, the set of classes in ImageNet-1k tend to be Anglo-centric due to inherited biases from WordNet (Shankar et al., 2017; DeVries et al., 2019; Liu et al., 2021) so while our benchmark evaluates the performance on languages from all over the globe, we do not evaluate the model performance on concepts specific (or even unique) to cultures in which the languages are spoken. As a result, Babel-ImageNet may overestimate the actual usability of an VL model for real-world uses in some cultures and geographies. Further, we select for Babel-ImageNet the 92 languages used in XLM-R pretraining as a tradeoff between language coverage and usability. This selection reinforces research focus on those languages to the detriment of other (mainly extremely low-resource) languages. However, we release our code, as well as data for labels of 298 languages and encourage future research to consider an even wider set of languages.
Appendix C Data and Training Details
C.1 Babel-ImageNet
af 303 am 85 ar 636 as 98 az 365 be 415 bg 602 bn 282 br 297
bs 156 ca 767 cs 615 cy 407 da 610 de 738 el 572 eo 603 es 845
et 496 eu 625 fa 682 fi 973 fr 799 fy 155 ga 502 gd 217 gl 473
gu 106 ha 47 he 648 hi 342 hr 347 hu 594 hy 454 id 463 is 409
it 773 ja 733 jv 183 ka 438 kk 365 km 167 kn 175 ko 648 ku 101
ky 247 la 276 lo 141 lt 535 lv 392 mg 64 mk 453 ml 281 mn 201
mr 140 ms 419 my 232 ne 134 nl 749 no 599 om 18 or 71 pa 128
pl 778 ps 112 pt 667 ro 687 ru 748 sa 66 sd 61 si 97 sk 509
sl 393 so 58 sq 273 sr 468 su 98 sv 699 sw 220 ta 346 te 202
th 896 tl 272 tr 559 ug 106 uk 640 ur 220 uz 254 vi 523 xh 35
yi 175 zh 885
@@ C.1 Babel-ImageNet
Table 4: The 92 languages of Babel-ImageNet in alphabetical order with the corresponding number of classes in Babel-ImageNet.
@@
Table 4 lists of the 92 Babel-ImageNet languages with their corresponding number of classes. Figure 5 visualizes the relationship between the number of classes of a language in Babel-ImageNet and the number of tokens for the language in the XLM-R pretraining corpus (which we use as a proxy for the language “resourceness”). We see that the two are generally correlated (Spearman rank correlation of 0.78), albeit with some expected outliers, e.g., Chinese is ”token-compact“ so the token count does not reflect its high-resourceness well.
C.2 Prompts
@@
We use NLLB (Costa-jussà et al., 2022) (nllb-200-distilled-1.3B) to translate the 80 prompts used by Radford et al. (2021) to our 92 languages. The exceptions are fy, la, and br, which are not supported by NLLB; for those languages we report the better results between: (1) using only the language-specific labels and (2) inserting ‘labels’ into English prompts. For the ISO 639-1 languages corresponding to macrolanguages, there is only one corresponding ISO 639-3 language in NLLB, except for no where we choose Bokmål and for az where we choose North Azerbaijani. We translate the prompts in their template form with the {} placeholders. We use a range of different methods like HTML tags or other special characters to increase the likelihood of preserving the placeholders during translation and then select the first successful approach. If no method worked, we append {} to the end of the sentence. We perform no language-specific adaptions like combining prompt variants with definite and indefinite articles for languages where this distinction does not exist (or articles do not exist at all) nor do we account for the grammatical gender of the classes when inserting them in the template.
C.3 Training
@@
Training Data: For the language-specific adaptation training in §6, we leveraged the BLIP (Li et al., 2022) image-caption dataset ccs_synthetic_filtered_large.json [^11] https://github.com/salesforce/BLIP#pre-training-datasets-download. Hyperparameters: We train with AdamW (Loshchilov and Hutter, 2019), 0.1 weight decay, a linear learning rate schedule with 20% warmup, learning rate 1e-3 (chosen with sweep over 1e-3, 5e-4, 3e-4, 1e-4), batch size 512 (OpenCLIP)/ 192 (M-CLIP), for 100 epochs (OpenCLIP)/ 15 epochs (M-CLIP; longer training yielded no improvements). Hyperparameters are chosen based on results on Sinhala. We perform no early stopping and use the last epoch for evaluation. The temperature for the contrastive loss is a trainable parameter as in Radford et al. (2021) but we freeze it for text contrastive loss (training it resulted in worse results). The maximum text sequence length is 70. For adapters, we use the Pfeiffer architecture (Pfeiffer et al., 2020b) (task adapters, not language adapters) with reduction factor 16 with the implementation from AdapterHub (Pfeiffer et al., 2020a). We pre-encode images and English captions; i.e. the English embeddings for MSE and contrastive loss are not computed by the trained model but come from the model before training. We do not use any type of image augmentation. Negative results: We experimented with the following methods but did not pursue them further due to not-better or poor results.
@@
Appendix D Further Experiments and Analysis
D.1 Experimental Validation of Machine-translated Prompts
@@
D.2 Comparison with Existing ImageNet Translations
@@
D.3 Performance Differences between Distilled and Not-Distilled Languages
@@
D.4 LAION5B: Language Distribution and Performance
@@
The distillation-based models evaluated in our benchmark use the same number of training examples for every language. The OpenCLIP models, on the other hand, are trained on LAION5B which follows a more ‘natural’ distribution of image-caption pairs across languages, as found on the web: Figure 7(b) shows that over half the data is English, 7 high-resource languages account for another 25% of the data, whereas all remaining languages “share” the remaining 25%. We can see in Figure 7(a) that the OpenCLIP’s ZS-IC performance for Babel-ImageNet languages highly depends on the number of instances of those languages in the LAION5B dataset. The Spearman rank correlation between the number of language-specific LAION5B examples and the respective Babel-ImageNet accuracy for the language is 0.76. This suggests that pure image-text contrastive pre-training results in poor generalization and limited cross-lingual gains to languages unseen in pretraining. Additional training objectives that aim to better align the multilingual space for example using paired text like in MURAL (Jain et al., 2021) might be necessary to improve results of OpenCLIP-like models (trained from scratch) for low-resource languages.
Appendix E Full Results

### v2

@@ 7 Conclusion
We introduced Babel-ImageNet, the first massively multilingual translation of the ImageNet classes to 100 languages. We leverage the WordNet synsets as the link between ImageNet and BabelNet to obtain high-quality translations without relying on MT or human annotators. Using Babel-ImageNet, we carried out the most comprehensively multilingual comparative evaluation of 11 public CLIP models on zero-shot image classification, demonstrating that all models fail for low(er)-resource languages. Crucially, we validate our benchmark by showing that models’ text-to-image retrieval performance (on three datasets) strongly correlates with their ZS-IC performance on Babel-ImageNet for the corresponding languages. Finally, we proposed a parameter-efficient fine-tuning procedure that drastically improves the performance of multilingual CLIP models for low-resource languages.
@@
8 Limitations
While Babel-ImageNet greatly improves language coverage for evaluation of multilingual VL models, there are some limitations of our work:
@@ 8 Limitations
For one, ImageNet classes tend to be Anglo-centric due to inherited biases from WordNet (Shankar et al., 2017; DeVries et al., 2019; Liu et al., 2021) so while our benchmark evaluates the performance on languages from all over the globe, we do not evaluate the model performance on concepts specific (or even unique) to cultures in which those languages are spoken. As a result, Babel-ImageNet may overestimate the actual usability of an VL model for real-world uses in some cultures and geographies.
Further, we select for Babel-ImageNet the 92 languages used in XLM-R pretraining along with 8 more manually chosen languages. This selection reinforces research focus on those languages to the detriment of other (mainly extremely low-resource) languages. However, we release our code, as well as data for labels of 298 languages and encourage future research to consider an even wider set of languages.
Acknowledgements
This work was in part supported by the Alexander von Humboldt Foundation.
Appendix A License
Babel-ImageNet is a processed version of BabelNet v5.2 downloaded from https://babelnet.org, made available with the BabelNet Non-Commercial License (see https://babelnet.org/full-license).
Appendix B Data and Training Details
B.1 Babel-ImageNet
very low om (18), xh (35), ha (47), so (58), sd (61), nah (62), hak (63), mg (64), sa (66), or (71), ce (73), chr (83), am (85), diq (89), si (97), su (98), as (98)
low ku (101), gu (106), ug (106), ps (112), pa (128), ne (134), cv (137), mr (140), lo (141), fy (155), bs (156), km (167), kn (175), yi (175), jv (183), mn (201), te (202), gd (217), sw (220), ur (220), my (232), ky (247), uz (254), nv (257), tl (272), sq (273), la (276), wuu (278), ml (281), bn (282), br (297), af (303)
mid hi (342), ta (346), hr (347), az (365), kk (365), lv (392), sl (393), cy (407), is (409), be (415), ms (419), ka (438), mk (453), hy (454), id (463), sr (468), gl (473), et (496), ga (502), sk (509), vi (523), lt (535), tr (559), el (572), hu (594), no (599), bg (602), eo (603), da (610), cs (615), eu (625), ar (636), uk (640), ko (648), he (648)
high pt (667), fa (682), ro (687), sv (699), ja (733), de (738), ru (748), nl (749), ca (767), it (773), pl (778), fr (799), es (845), zh (885), th (896), fi (973)
@@ B.1 Babel-ImageNet
Table 5: The 100 languages of Babel-ImageNet in their respective groups. Number of classes in parentheses.
@@
Table 5 lists the 100 Babel-ImageNet languages with their corresponding number of classes.
@@
Figure 4 visualizes the relationship between the number of classes of a language in Babel-ImageNet and the number of tokens for the language in the XLM-R pretraining corpus (which we use as a proxy for the language “resourceness”). We see that the two are generally correlated (Spearman rank correlation of 0.78), albeit with some expected outliers, e.g., Chinese is ”token-compact“ so the token count does not reflect its high-resourceness well.
B.2 Prompts
@@
We use NLLB (Costa-jussà et al., 2022) (nllb-200-distilled-1.3B) to translate the 80 prompts used by Radford et al. (2021) to our 100 languages. The exceptions are fy, la, br, wuu, nv, cv, diq, chr, ce, hak, nah, which are not supported by NLLB; for those languages we report the better results between: (1) using only the language-specific labels and (2) inserting ‘labels’ into English prompts. For the ISO 639-1 languages corresponding to macro languages, there is only one corresponding ISO 639-3 language in NLLB, except for no where we choose Bokmål and for az where we choose North Azerbaijani. We translate the prompts in their template form with the {} placeholders. We use a range of different methods like HTML tags or other special characters to increase the likelihood of preserving the placeholders during translation and then select the first successful approach. If no method worked, we append {} to the end of the sentence. We perform no language-specific adaptions like combining prompt variants with definite and indefinite articles for languages where this distinction does not exist (or articles do not exist at all) nor do we account for the grammatical gender of the classes when inserting them in the template.
B.3 Training
Training Data: For the language-specific adaptation training in §6, we leveraged the BLIP (Li et al., 2022) image-caption dataset ccs_synthetic_filtered_large.json [^9] https://github.com/salesforce/BLIP#pre-training-datasets-download.
@@ B.3 Training
Hyperparameters: We train with AdamW (Loshchilov and Hutter, 2019), 0.1 weight decay, a linear learning rate schedule with 20% warmup, learning rate 1e-3 (chosen with sweep over 1e-3, 5e-4, 3e-4, 1e-4), batch size 512 (OpenCLIP)/ 192 (M-CLIP), for 100 epochs (OpenCLIP)/ 15 epochs (M-CLIP; longer training yielded no improvements). Hyperparameters are chosen based on results on Sinhala. We perform no early stopping and use the last epoch for evaluation. The temperature for the contrastive loss is a trainable parameter as in Radford et al. (2021) but we freeze it for text contrastive loss (training it resulted in worse results). The maximum text sequence length is 70. For adapters, we use the Pfeiffer architecture (Pfeiffer et al., 2020b) (task adapters, not language adapters) with reduction factor 16 with the implementation from AdapterHub (Pfeiffer et al., 2020a). We pre-encode images and English captions; i.e. the English embeddings for MSE and contrastive loss are not computed by the trained model but come from the model before training. We do not use any type of image augmentation.
Negative results: We experimented with the following methods but did not pursue them further due to not-better or poor results.
@@
Appendix C Further Experiments and Analysis
C.1 Experimental Validation of Machine-translated Prompts
@@
C.2 Comparison with Existing ImageNet Translations
@@
C.3 Performance Differences between Distilled and Not-Distilled Languages
@@
C.4 LAION5B: Language Distribution and Performance
@@
The distillation-based models evaluated in our benchmark use the same number of training examples for every language. The OpenCLIP models, on the other hand, are trained on LAION5B which follows a more ‘natural’ distribution of image-caption pairs across languages, as found on the web: Figure 7(b) shows that over half the data is English, 7 high-resource languages account for another 25% of the data, whereas all remaining languages “share” the remaining 25%.
We can see in Figure 7(a) that the OpenCLIP’s ZS-IC performance for Babel-ImageNet languages highly depends on the number of instances of those languages in the LAION5B dataset. The Spearman rank correlation between the number of language-specific LAION5B examples and the respective Babel-ImageNet accuracy for the language is 0.76. This suggests that pure image-text contrastive pre-training results in poor generalization and limited cross-lingual gains to languages unseen in pretraining. Additional training objectives that aim to better align the multilingual space for example using paired text like in MURAL (Jain et al., 2021) might be necessary to improve results of OpenCLIP-like models (trained from scratch) for low-resource languages.
Appendix D Full Results
@@
lang AC L-14 MC mB B-32 MC B-16+ MC B-32 MC L-14 NS-base NS-large OAI B-32 OC B-32 OC H-14 mSigLIP ST B-32
en 73.36 29.97 47.02 44.06 52.34 39.75 51.96 63.35 62.32 76.95 75.12 39.45
af 22.6 25.26 50.33 47.48 55.32 41.11 53.89 10.65 36.48 47.53 47.17 14.94
am 7.13 0.99 27.22 29.25 29.11 50.12 55.79 1.51 1.04 2.56 14.56 0.85
ar 42.44 13.99 32.06 30.68 36.3 29.52 40.1 0.56 24.51 31.7 42.27 10.69
as 6.53 17.22 21.69 22.84 24.92 39.94 49.22 1.18 5.43 9.53 23.14 2.92
az 16.4 23.71 26.57 25.69 28.81 34.77 44.24 7.27 25.83 36.98 47.82 12.68
be 27.52 14.08 35.35 32.66 36.85 36.72 48.4 0.64 27.18 40.12 32.05 10.02
bg 34.48 21.61 42.63 40.57 47.13 32.23 42.43 0.8 41.26 54.02 58.99 19.56
bn 7.25 25.99 31.17 29.77 35.55 47.48 57.68 0.17 8.61 19.57 49.42 3.06
br 12.02 3.99 10.71 9.64 11.61 8.82 12.28 7.14 13.02 15.23 14.96 3.47
bs 29.45 39.15 60.18 59.74 65.27 49.55 58.69 13.04 55.99 69.88 79.49 35.35
ca 32.02 19.9 37.48 34.94 40.33 28.34 38.67 11.54 33.37 46.78 42.96 17.5
ce 16.38 6.9 15.73 11.37 13.53 19.04 20.22 1.18 21.21 22.19 25.53 4.36
chr 1.2 1.2 1.2 1.16 1.47 1.18 2.19 1.47 1.33 0.48 1.54 1.64
cs 20.03 22.04 38.44 36.59 41.89 31.26 42.06 6.36 41.8 54.11 66.02 19.5
cv 29.93 11.75 23.64 19.78 22.99 17.55 24.67 0.58 31.8 34.67 32.96 10.32
cy 12.8 19.81 11.39 11.31 12.99 25.8 33.45 5.04 12.08 16.52 12.72 5.62
da 24.96 22.79 46.92 44.45 51.84 34.66 46.87 12.39 42.92 55.5 62.46 22.65
de 27.96 20.07 42.01 39.59 46.39 31.74 44.43 15.83 47.58 61.05 67.1 18.7
diq 14.81 6.49 20.67 19.19 18.47 17.62 22.11 8.2 18.72 21.42 25.96 9.08
el 5.03 16.58 41.73 39.03 45.75 33.3 45.72 0.9 37.21 50.77 51.16 15.29
eo 24.84 13.04 25.22 24.64 28.58 35.12 48.63 5.77 21.82 28.5 21.65 10.9
es 51.78 20.13 38.64 36.31 42.07 32.14 42.42 16.98 44.78 57.4 60.72 20.67
et 15.54 19.79 45.19 42.22 48.73 29.21 39.91 4.64 29.68 37.84 53.5 13.51
eu 17.72 9.63 18.92 18.72 21.25 34.16 45.22 7.43 16.87 21.51 21.89 9.0
fa 20.13 20.69 28.24 25.55 29.26 27.6 37.34 0.46 32.66 42.48 52.3 13.89
fi 10.56 14.45 27.49 25.79 30.27 22.82 32.19 3.96 25.0 35.69 46.88 11.25
fr 54.06 20.17 38.84 35.84 42.39 32.21 43.05 21.08 46.21 59.72 63.54 20.93
fy 23.73 6.83 27.9 28.72 30.09 22.7 29.64 11.99 34.13 40.65 36.21 6.15
ga 8.36 3.12 9.42 8.29 9.76 20.12 27.98 2.72 6.71 9.45 5.55 2.24
gd 6.01 3.46 8.25 7.52 10.53 25.43 30.48 3.05 5.98 7.3 4.94 2.63
gl 43.92 23.08 41.53 40.26 45.05 35.13 47.57 15.47 43.82 54.66 47.37 23.91
gu 11.34 28.77 28.11 26.94 32.32 49.3 63.42 0.4 6.94 12.02 20.72 7.64
ha 3.23 19.36 4.94 4.17 5.19 41.96 49.74 1.62 2.47 3.57 3.06 2.0
hak 33.27 18.7 32.29 26.03 28.1 21.94 27.9 3.97 32.6 35.08 31.14 11.08
he 7.03 16.68 22.04 21.02 23.7 26.07 35.16 0.3 26.64 35.87 48.99 12.86
hi 12.89 20.99 37.37 37.04 41.61 41.47 54.75 0.09 25.85 38.65 40.51 17.58
hr 21.75 30.22 51.08 50.67 57.0 40.19 51.75 8.63 44.82 58.2 66.58 25.64
hu 16.55 19.91 45.47 42.69 50.3 31.38 42.39 5.57 40.93 53.25 63.9 17.1
hy 5.3 16.89 18.65 17.97 18.26 26.7 36.7 0.17 9.91 17.5 35.09 7.43
id 26.52 26.6 50.39 47.73 55.14 42.63 53.63 16.09 43.78 57.61 69.25 23.49
is 10.82 18.91 41.84 39.96 45.56 29.15 38.46 2.93 13.51 19.25 30.18 6.04
it 49.43 19.97 37.85 35.81 41.27 29.95 41.32 15.55 41.88 55.03 59.89 19.49
ja 56.33 18.55 25.57 24.93 29.2 28.91 37.71 4.18 42.35 57.14 62.79 15.74
jv 21.8 19.51 33.85 32.89 37.27 42.74 52.2 14.89 31.81 40.44 50.4 15.08
ka 9.0 15.48 22.83 22.05 24.26 26.69 34.27 0.2 11.09 20.44 41.13 8.66
kk 31.61 19.96 26.59 25.59 28.44 34.44 44.46 0.53 28.12 34.52 47.7 10.8
km 6.72 0.91 16.57 17.68 19.77 25.78 30.16 0.79 3.31 3.16 31.54 0.38
kn 10.97 21.17 27.09 27.36 29.57 43.3 53.95 1.04 4.16 5.68 16.3 2.16
ko 53.8 16.02 24.02 21.61 25.12 24.48 33.15 0.42 33.52 43.51 56.58 12.52
ku 11.23 8.95 12.75 14.2 14.79 14.73 21.27 5.09 14.1 15.88 15.5 8.73
ky 32.11 18.48 28.79 28.7 31.69 38.53 47.82 0.65 32.56 38.55 41.6 14.01
la 23.06 4.7 16.43 11.23 14.38 12.8 21.94 10.77 21.14 26.65 26.35 4.2
lo 9.86 0.74 12.45 12.75 13.05 32.13 37.59 0.77 0.85 2.21 10.26 0.71
lt 18.7 21.69 29.9 28.06 31.96 29.09 39.78 5.94 34.68 45.79 55.12 18.58
lv 21.42 26.01 33.9 34.14 37.14 31.99 40.98 7.74 38.52 47.3 56.7 23.65
mg 13.25 8.62 13.06 14.19 14.12 38.47 47.72 4.06 13.94 14.62 12.34 9.06
mk 31.41 23.55 47.78 46.53 51.4 33.18 43.89 1.33 36.08 49.47 47.31 20.45
ml 7.64 20.05 38.69 38.8 44.01 37.69 49.56 0.16 1.69 4.45 15.31 1.23
mn 20.44 28.29 20.86 19.37 21.78 25.09 28.6 1.02 18.58 26.49 44.96 14.06
mr 17.46 25.9 45.89 45.41 47.83 59.39 68.63 1.13 32.29 41.66 42.24 29.66
ms 21.25 22.12 39.23 37.86 43.59 35.71 46.72 14.94 37.08 48.35 61.1 19.26
my 9.47 2.49 14.72 14.49 16.2 38.34 46.23 0.49 5.82 10.27 17.16 10.35
nah 8.13 5.52 6.52 7.42 9.03 7.13 7.97 4.35 10.68 11.35 9.9 2.23
ne 16.51 17.07 37.57 40.16 39.21 49.78 60.39 0.54 25.0 36.39 40.16 16.9
nl 25.03 18.57 39.09 37.33 44.56 29.67 41.41 13.05 41.71 54.93 57.96 18.88
no 23.47 21.13 43.4 40.66 47.56 33.24 43.78 9.4 41.98 53.57 61.21 19.55
nv 0.65 0.39 0.6 0.38 0.16 0.66 0.29 0.35 0.44 0.41 0.25 0.54
om 8.0 3.44 11.11 9.67 9.67 29.22 30.44 4.44 5.44 13.0 10.22 8.78
or 12.62 1.75 30.62 31.18 36.11 57.24 70.56 1.75 3.27 1.61 1.61 1.77
Table 8: Babel-ImageNet results for all languages (sorted alphabetically expect for English). To save space, we shorten sources (OAI: OpenAI; OC: OpenCLIP; MC: M-CLIP; ST: SentenceTransformer, AC: AltCLIP, NS: NLLB-SigLIP) and remove the text model if possible.
lang AC L-14 MC mB B-32 MC B-16+ MC B-32 MC L-14 NS-base NS-large OAI B-32 OC B-32 OC H-14 mSigLIP ST B-32
pa 11.42 7.97 33.89 32.09 31.69 60.27 70.47 1.53 1.77 2.56 3.05 4.09
pl 19.72 18.34 35.83 33.33 38.95 28.89 39.18 7.63 39.15 51.46 61.38 16.32
ps 18.05 10.86 19.48 18.98 23.21 33.38 41.8 1.57 21.66 25.04 24.52 11.88
pt 37.41 21.57 43.24 40.06 47.36 34.42 46.26 14.43 47.21 59.62 61.7 23.03
ro 23.92 18.78 40.45 37.78 44.26 29.74 41.21 10.19 35.17 47.6 51.22 16.22
ru 48.87 18.34 36.82 35.25 41.7 30.47 41.96 0.54 42.95 58.21 60.16 16.39
sa 11.18 9.45 17.36 17.21 18.97 32.36 37.09 0.91 11.55 11.73 23.15 9.55
sd 15.93 20.98 28.13 27.9 29.87 43.34 51.84 2.0 14.56 16.49 15.61 9.41
si 19.4 2.19 33.32 33.75 36.39 55.38 65.48 2.64 3.22 5.32 29.13 2.12
sk 22.97 25.31 43.22 40.56 46.11 35.56 46.68 8.65 44.99 58.68 66.17 22.72
sl 17.39 24.61 45.05 42.98 48.71 34.15 47.76 6.01 36.93 48.65 59.23 22.46
so 3.76 17.55 10.38 12.55 12.14 29.97 38.86 2.48 6.66 7.41 7.07 5.17
sq 21.71 26.21 49.82 48.12 54.07 38.39 48.67 8.24 33.99 43.35 43.77 24.86
sr 30.64 19.18 45.99 44.58 49.92 32.35 44.06 1.18 33.93 47.7 45.43 17.49
su 18.73 16.04 28.16 29.22 29.33 37.39 44.04 12.06 27.71 30.61 35.9 11.67
sv 21.06 21.1 45.72 42.0 49.44 31.93 42.26 9.17 42.88 55.2 61.67 19.68
sw 10.37 17.3 37.05 36.85 39.38 31.87 38.76 4.72 10.25 12.91 20.17 6.63
ta 5.53 18.67 18.78 18.09 20.98 30.98 41.14 0.35 4.54 6.73 30.21 2.12
te 10.75 24.95 32.56 31.75 34.44 45.62 58.13 0.51 2.83 3.71 19.17 3.02
th 11.85 15.55 29.69 27.77 32.77 25.63 33.35 1.33 28.71 40.16 40.84 10.62
tl 18.25 16.17 33.43 32.51 38.11 26.43 33.62 7.86 17.54 21.73 33.05 8.18
tr 17.67 23.16 44.98 42.78 48.4 35.84 46.72 7.61 41.25 52.88 63.65 19.47
ug 9.06 2.17 13.19 11.68 12.75 33.42 40.3 1.19 3.06 4.64 6.19 2.3
uk 34.82 17.61 36.67 35.77 41.22 28.99 39.5 0.63 39.19 53.18 57.54 15.13
ur 18.48 27.24 27.15 29.85 31.17 47.62 58.92 0.59 26.01 37.12 29.48 17.75
uz 17.24 21.4 22.57 22.54 24.13 36.09 48.54 6.13 21.83 28.45 36.42 9.8
vi 11.69 19.09 39.59 37.94 44.58 29.77 39.13 6.94 40.92 53.39 59.9 17.86
wuu 72.53 24.4 49.64 28.58 48.78 31.35 41.76 3.43 59.02 70.81 72.22 19.69
xh 21.77 16.34 19.2 17.71 20.11 57.6 69.2 18.86 24.11 27.14 24.8 14.23
yi 5.29 1.04 18.37 18.9 19.38 39.22 52.81 0.71 2.49 4.73 3.82 1.33
zh 53.35 21.99 36.41 33.5 40.9 25.5 33.26 1.85 40.73 53.28 55.42 19.75